# LLaVA-MORE [Uncaptioned image]: A Comparative Study of LLMs and Visual Backbones for Enhanced Visual Instruction Tuning

### v1

@@ 1 Introduction
Figure 1: Performance comparison of the best version of LLaVA-MORE with other LLaVA versions across different benchmarks for visual question answering and multimodal reasoning.
@@
The emergence of Large Language Models (LLMs) with remarkable expressive capabilities has revolutionized the way diverse language-related tasks are approached [16, 59, 62, 1]. This advancement has inspired the Computer Vision community to move beyond traditional text-only paradigms and adopt multiple modalities, including vision, audio, and beyond. Consequently, this shift has led to the emergence of Multimodal Large Language Models (MLLMs) [8], which establish sophisticated relationships between concepts across different embedding spaces, enabling richer multimodal understanding.
@@
Current MLLMs [5, 40, 41, 4, 15, 25] typically integrate a language model with a visual backbone using specialized adapters that bridge the gap between modalities. While these systems demonstrate impressive performance, the field has converged around a somewhat narrow technical approach, with most implementations leveraging LLaMA-derived language models and LLaVA-based training protocols. Additionally, visual encoders based on contrastive training such as CLIP [49] and its derivatives [64, 20, 69] have become the default choice for extracting visual features aligned with textual representations. These encoders are specifically trained to generate embeddings that seamlessly integrate with language models, further driving their widespread adoption. While contrastive learning has been highly effective in aligning images and text within a shared representation space, other vision models [9, 46] capable of capturing richer structural information have received comparatively less attention.
@@
To address this, our work conducts a comprehensive empirical study that systematically pairs diverse LLMs –ranging from efficient models [1] to significantly larger architectures [62, 59] – with various visual backbones [49, 69, 46, 64]. By exploring different architectural combinations, we aim to uncover the strengths and limitations of various vision-language integration strategies, shedding light on overlooked design choices and their impact on multimodal learning. Fig. 1 illustrates a comparison of our best-performing LLaVA-MORE models against various LLaVA-based competitors.
@@
To ensure experimental consistency, we follow the established LLaVA [41] methodology, pre-training models on natural language description tasks before applying visual instruction fine-tuning to improve cross-domain generalization and human alignment. However, recent works not only introduce architectural enhancements but also incorporate specially curated datasets [19, 7, 43], making fair comparisons across models challenging. To address this, we analyze performance variations across diverse datasets, using standardized LLaVA evaluations as a benchmark.
@@
To summarize, this work provides key insights into cross-modal representation learning and offers a practical guide for developing more efficient and effective MLLMs, while challenging conventional assumptions about dataset requirements for optimizing pre-training strategies.
@@ 2 Related Works
From that point onward, the development of more powerful MLLMs accelerated. Recent works have not only refined standard multimodal capabilities but also expanded their focus to a broader range of vision-centric benchmarks and diverse settings [7]. To achieve state-of-the-art results, some approaches prioritize enhancing the visual encoding phase, eventually training a vision encoder from scratch [4]. Others focus instead on scaling up spatial and temporal resolutions or improving visual token compression [43].
Despite the impressive performance of these models, they rely on specially curated datasets during training [30, 19, 7, 43, 31], which complicates fair evaluation and direct comparisons across different architectures and settings.
Large Language Models and Recent Advances. Despite the development of various LLM architectures [70, 65, 50, 17, 67], the LLaMA family [62, 63, 22] remains a popular choice thanks to its open-access nature, exclusive reliance on publicly available training data, and availability in multiple model sizes. Variants such as Alpaca [56] and Vicuna [16] further build upon LLaMA, enhancing its instruction-following and conversational abilities.
@@
LLaMA-3 [22] marks a major breakthrough in LLMs, offering 8B to 405B parameter models with enhanced multilingual, coding, and reasoning abilities. Its improvements stem from rigorous data curation, pre-processing, and filtering, highlighting the critical role of high-quality training data in boosting efficiency and accuracy. This principle is exemplified by the Phi models [2, 35, 1], which outperforms larger LLMs across multiple benchmarks. Phi-4 [1] advances small-scale models with synthetic data, optimized training, and post-training techniques, showing that data refinement can rival model scaling.
Visual Backbones. Beyond advancements in language modeling, the visual backbone plays a crucial role in the performance of MLLMs by extracting and encoding image features for multimodal reasoning and understanding. The most widely used visual encoders are Vision Transformers (ViTs) trained with CLIP-style contrastive learning, with CLIP ViT-L [49] being the most commonly adopted architecture. Building on the success of CLIP, particularly its contrastive training approach, many studies explore similar methods leveraging cross-modal alignment.
@@ 3 Overall Architecture
Following the standard LLaVA architecture, a typical MLLM consists of three fundamental components: a large language model backbone for user interaction and generating text, one or more visual encoders to process visual input and extract features, and at least one vision-to-language adapter to bridge the gap between visual and textual modalities [8]. The visual encoder provides essential visual features to the LLM, with CLIP-based architectures [49, 66] being the most commonly used. In this setup, the visual encoder is typically pre-trained and kept frozen for the entire training of the MLLM architecture.
In this paper, we proposed LLaVA-MORE, a new family of models that extend the standard LLaVA architecture by combining the visual encoder with various LLMs, ranging from small- to medium-scale models. As small-scale models, we utilize Gemma-2 2B [59] and Phi-4-Mini [1] (with 3.8B parameters), both designed for strong reasoning scalability, effectively challenging larger models. For medium-scale models, we select the variant of Gemma-2 [59] with 9B parameters alongside two recent architectures: the original LLaMA-3.1 [22] LLM with 8B parameters and its distilled DeepSeek-R1 version [23] (i.e., DeepSeek-R1-Distill-LLaMA-8B). For each LLM category, we assess the impact of varying the visual backbone, aiming to identify the optimal configuration. Specifically, our study compares the standard CLIP ViT-L/14 employed in the original LLaVA-1.5 model [41] with two variants of DINOv2 with and without registers [46, 18], known for their strong, semantically rich features, as well as SigLIP [69] and its more advanced successor, SigLIP2 [64].
@@ 4.1 Implementation Details
Architectural Components. The LLaVA-MORE family follows the standard LLaVA architecture, employing CLIP ViT-L/14@336 as the visual backbone while varying the underlying language model. We categorize the selected LLMs into two groups based on scale: small-scale models, including Gemma-2-2B and Phi-4-3.8B, and medium-scale models, such as LLaMA-3.1-8B, DeepSeek-R1-Distill-LLaMA-8B, and Gemma-2-9B. To further investigate the impact of the visual backbone, we conduct experiments on the best-performing models, replacing CLIP [49] with alternative vision encoders, including DINOv2 [46], SigLIP [69], and SigLIP2 [64]. Additionally, we examine the effects of applying Scaling on Scales (S2[53] to both the CLIP and SigLIP2 architectures. Finally, motivated by the LLaVA-1.5 framework and insights from [11, 12], which highlight the advantages of using multilayer perceptrons over linear projections in self-supervised learning, we adopt a two-layer MLP as the vision-language adapter to enhance multimodal fusion.
Training Details. Following the LLaVA framework, we adopt a two-stage training strategy. In the first stage, only the weights of the adapter are updated using image-caption pairs as training data. Specifically, the caption style follows the alt-text structure as in web-scale multimodal datasets. The models are trained for one full epoch, covering a total of 558k samples from a combination of different sources (i.e., LAION [52], CC3M [10], and SBU [47]). In the second step, we fine-tune the model on high-quality visual instruction-following data to improve its multimodal reasoning capabilities. This sequential training approach has been shown to significantly improve performance on downstream tasks [40]. Notably, the next token prediction is used as the loss function in both training phases.
LLaVA-MORE models are trained with the same set of hyperparameters as LLaVA-1.5 to ensure consistency and comparability. In particular, we employ a global batch size of 256 during pre-training and 128 during the visual instruction tuning phase. All experiments are run in a multi-GPU, multi-node configuration with a total of 16 A100 64GB NVIDIA GPUs.
@@ 4.2 Evaluation Benchmarks
  • ScienceQA [45] evaluates models with challenging multimodal multiple-choice questions across three subjects (i.e., natural science, language science, and social science), 26 topics, 127 categories, and 379 skills. Each question is annotated with explanations linked to relevant lectures. We report results on the test set which includes 4,241 examples.
@@
  • POPE [36] benchmarks object hallucinations in MLLM generations, including several subsets, namely random, popular, and adversarial, which are generated using various sampling methodologies. It consists of 8,910 binary classification queries, enabling thorough analysis of the object hallucination phenomena in MLLMs.
@@
  • MMBench (MMB) [42] includes approximately 3,000 multiple-choice questions, distributed across 20 distinct domains. Questions are designed to assess the effectiveness of MLLMs across various task paradigms. These capabilities are systematically structured into a hierarchical taxonomy, encompassing broad categories like perception and reasoning, as well as finer-grained skills such as object localization and attribute inference.
@@
  • MMMU [68] is a challenging benchmark focusing on domain-specific knowledge of multimodal models. It consists of 900 validation samples drawn from university textbooks or online courses spanning six main disciplines. Questions may include multiple images interleaved with text. Evaluation includes exact and word matching for multiple-choice and open-ended questions, respectively, and models are tested in zero or few-shot settings.
@@ 4.3 Assessing the Optimal LLM Choice
Conversely, among medium-scale models, LLaVA-1.5-7B and LLaVA-1.5-LLaMA3-8B provide stronger baselines, with LLaVA-1.5-LLaMA3-8B achieving the highest score in MME-P (1544.4). However, our LLaVA-MORE models consistently surpass these baselines, demonstrating superior performance across both VQA and MLLM benchmarks. In particular, Gemma-2-9B emerges as the best-performing model, especially excelling in VQA benchmarks. It achieves the highest scores on GQA and AI2D, significantly outperforming both baselines and other models within the LLaVA-MORE family. In MLLM benchmarks, LLaVA-MORE combined with LLaMA-3.1-8B demonstrates instead strong capabilities on the MMB dataset. This highlights the model’s proficiency in understanding complex multimodal inputs and its ability to reason across different modalities. Notably, our models exhibit less sensitivity to object hallucinations, as seen by the results achieved on the POPE benchmark, and demonstrate superior performance on the SEED dataset, further highlighting their robustness in multimodal reasoning tasks.
@@
This trend extends to multimodal LLM benchmarks, where LLaVA-MORE models demonstrate competitive performance in MMB and SEED, further highlighting their efficiency and strong reasoning capabilities despite their smaller size. Overall, our results emphasize that scaling up is not the only path to better performance, as architectural choices and fine-tuning strategies significantly impact model effectiveness across different benchmarks.
@@ 4.4 Changing the Visual Backbone
We then investigate which visual backbone is more promising for building MLLMs. To this end, we select the best small- and medium-scale models resulting from Table 1 and study how their performance is affected when the visual encoder is varied. In detail, we opt for Phi-4-3.8B as the small-scale LLM (i.e., LLaVA-MORE-3.8B), and Gemma-2-9B as the medium-scale one (i.e., LLaVA-MORE-9B). As per the visual backbone, in Table 2, we include four pre-trained ViT-based models, in addition to the standard CLIP used by LLaVA. All models share the ViT-L/14 architecture, and yet there are striking differences in terms of the resolutions of the input image, and on the pre-training strategies. In particular, we can delineate two pre-training paradigms: DINOv2 [46], eventually enhanced with registers [18], has been pre-trained with self-supervision and knowledge distillation on 142M images, while the other visual encoders leverage the weak supervision of noisy image-text pairs during pre-training. Specifically, CLIP [49] learns robust visual models via contrastive learning, while SigLIP [69] exploits the sigmoid loss to improve cross-modal alignment. The recent SigLIP2 [64] builds upon SigLIP by incorporating additional training objectives, including image captioning, self-distillation, and image-masked prediction.

### v2

@@ 1 Introduction
Figure 1: Performance comparison of the best version of LLaVA-MORE with other LLaVA variants across different benchmarks for multimodal reasoning and visual question answering.
@@
The emergence of Large Language Models (LLMs) with remarkable expressive capabilities has revolutionized the way diverse language-related tasks are approached [17, 61, 64, 1]. This advancement has inspired the Computer Vision and Multimedia communities to move beyond traditional text-only paradigms and adopt multiple modalities, including vision, audio, and beyond. Consequently, this shift has led to the emergence of Multimodal Large Language Models (MLLMs) [9], which establish sophisticated relationships between concepts across different embedding spaces, enabling richer multimodal understanding.
@@
Current MLLMs [5, 41, 42, 4, 16, 26] typically integrate a language model with a visual backbone using specialized adapters that bridge the gap between modalities. While these systems demonstrate impressive performance, the field has converged around a somewhat narrow technical approach, with most implementations leveraging LLaMA-derived LLMs and LLaVA-based training protocols. Additionally, visual encoders based on contrastive training such as CLIP [51] and its derivatives [66, 21, 71] have become the default choice for extracting visual features. These encoders are specifically trained to generate embeddings that seamlessly integrate with language models, further driving their widespread adoption. While contrastive learning has been highly effective in aligning images and text within a shared space, other vision models [10, 48] learn robust visual features in a purely self-supervised scheme, without relying on weak supervision from text. These models showcase intriguing emerging properties, and yet their application to MLLMs is relatively understudied.
@@
To address this, our work conducts a comprehensive empirical study that systematically pairs diverse LLMs –ranging from efficient models [1] to significantly larger architectures [64, 61] – with various visual backbones [51, 71, 48, 66]. By exploring different architectural combinations, we aim to uncover the strengths and limitations of various vision-language integration strategies, shedding light on overlooked design choices and their impact on multimodal learning. Fig. 1 illustrates a comparison of our best-performing model (i.e., LLaVA-MORE-9B, based on SigLIP2 as visual encoder and Gemma-2-9B as LLM) against LLaVA-based competitors (i.e., LLaVA-1.5-7B [42] and LLaVA-1.5-LLaMA3-8B [53], both using a CLIP-based visual encoder and Vicuna-7B and LLaMA3-8B as LLM, respectively).
@@
To ensure experimental consistency, we follow the established LLaVA [42] methodology, pre-training models on natural language description tasks before applying visual instruction fine-tuning to improve cross-domain generalization and human alignment. However, recent works not only introduce architectural enhancements but also incorporate specially curated datasets [20, 7, 44], making fair comparisons across models challenging. To isolate the role of data, in this work we compare LLaVA models pre-trained on different datasets against the same evaluation protocol.
@@
To summarize, this work provides key insights into cross-modal representation learning and offers a practical guide for developing more efficient and effective MLLMs, while challenging conventional assumptions about visual and textual backbones and dataset requirements for optimizing pre-training strategies.
@@ 2 Related Works
Recent works have not only refined standard multimodal capabilities but also expanded their focus to a broader range of vision-centric benchmarks and diverse settings [7]. To achieve state-of-the-art results, some approaches prioritize enhancing the visual encoding phase, eventually fine-tuning [47] a vision encoder or even training it from scratch [4]. Others focus instead on scaling up spatial and temporal resolutions or improving visual token compression [44, 8] for better efficiency.
Despite the impressive performance of these models, a significant drawback lies in their reliance on specially curated datasets during training [31, 20, 7, 44, 32], which complicates fair evaluation and direct comparisons across different architectures and settings.
Large Language Models and Recent Advances. Despite the development of various LLM architectures [72, 67, 52, 18, 69], the LLaMA family [64, 65, 23] remains a popular choice thanks to its open-access nature, reliance on publicly available model weights, and availability in multiple sizes. Variants such as Alpaca [58] and Vicuna [17] further build upon LLaMA, enhancing its instruction-following and conversational abilities.
@@
A major breakthrough in LLM development is the introduction of LLaMA-3 [23], offering 8B to 405B parameter models with enhanced multilingual, coding, and reasoning abilities. Its improvements stem from rigorous data curation, pre-processing, and filtering, highlighting the critical role of high-quality training data in boosting efficiency and accuracy. This principle is further exemplified by the Phi models [2, 36, 1], which outperforms larger LLMs across multiple benchmarks. Among these, Phi-4 [1] pushes the boundaries of small-scale models through synthetic data generation, optimized training, and advanced post-training techniques, demonstrating that strategic data refinement can rival sheer model scaling.
Visual Backbones. Beyond advancements in language modeling, the visual backbone plays a crucial role in the performance of MLLMs by extracting and encoding image features for multimodal reasoning and understanding. The most widely used visual encoders are Vision Transformers (ViTs) trained with CLIP-style contrastive learning, with CLIP ViT-L [51] being the most commonly adopted architecture. Building on the success of CLIP, several studies have explored similar methodologies that leverage the inherent alignment of cross-modal embeddings.
@@ 3 Overall Architecture
Following the LLaVA architecture, a typical MLLM consists of three fundamental components: a large language model backbone for user interaction and generating text, one or more visual encoders to process visual input and extract features, and at least one vision-to-language adapter to bridge the gap between visual and textual modalities [9]. The visual encoder provides essential global visual features to the LLM, with CLIP-based architectures [51, 68] being the most commonly used. In this setup, the visual encoder is typically pre-trained on another task and kept frozen for the entire training of the MLLM.
In this paper, we propose LLaVA-MORE, a new family of models that extend the standard LLaVA architecture by combining the visual encoder with various LLMs, ranging from small- to medium-scale models. As small-scale models, we utilize Gemma-2 2B [61] and Phi-4-Mini [1] (with 3.8B parameters), both designed for strong reasoning scalability, effectively challenging larger models. For medium-scale models, we select the variant of Gemma-2 [61] with 9B parameters alongside two recent architectures: the original LLaMA-3.1 [23] LLM with 8B parameters and its distilled DeepSeek-R1 version [24] (i.e., DeepSeek-R1-Distill-LLaMA-8B). For each LLM category, we assess the impact of varying the visual backbone, with the aim to identify the optimal configuration. Specifically, our study compares the standard CLIP ViT-L/14 encoder employed in the LLaVA-1.5 model [42] with two variants of DINOv2 differentiated by the presence or the absence of visual register tokens [48, 19], known for their strong, semantically rich visual features, as well as SigLIP [71] and its more advanced successor, SigLIP2 [66].
@@ 4.1 Implementation Details
Architectural Components. The LLaVA-MORE family follows LLaVA architecture, employing CLIP ViT-L/14@336 as the visual backbone while varying the underlying language model. We categorize the selected LLMs into two groups based on scale: small-scale models, including Gemma-2-2B and Phi-4-3.8B, and medium-scale models, such as LLaMA-3.1-8B, DeepSeek-R1-Distill-LLaMA-8B, and Gemma-2-9B. To further investigate the impact of the visual backbone, we conduct experiments on the best-performing models, replacing CLIP [51] with alternative vision encoders, including DINOv2 [48], SigLIP [71], and SigLIP2 [66]. Additionally, we examine the effects of applying Scaling on Scales (S2[55] to both the CLIP and SigLIP2 architectures. Finally, motivated by the LLaVA-1.5 framework and insights from [12, 13], which highlight the advantages of using MLPs over linear projections in self-supervised learning, we adopt a two-layer MLP as the vision-language adapter to enhance multimodal fusion.
Training Details. Considering the LLaVA framework, we adopt a two-stage training strategy. In the first stage, only the weights of the adapter are updated using image-caption pairs as training data. Specifically, the caption style follows the alt-text structure as in web-scale multimodal datasets. The models are trained for one full epoch, covering a total of 558k samples from a combination of different sources (i.e., LAION [54], CC3M [11], and SBU [49]). In the second step, we fine-tune the model on high-quality visual instruction-following data to improve its multimodal reasoning capabilities. This sequential training approach has been shown to significantly improve performance on downstream tasks [41]. Notably, the next token prediction is used as the loss function in both training phases. LLaVA-MORE models are trained with the same set of hyperparameters as LLaVA-1.5 to ensure consistency and comparability. In particular, we employ a global batch size of 256 during pre-training and 128 during the visual instruction tuning phase. All experiments are run in a multi-GPU, multi-node configuration with a total of 16 A100 64GB NVIDIA GPUs.
@@ 4.2 Evaluation Benchmarks
  • ScienceQA [46] evaluates models with challenging multimodal multiple-choice questions across three diverse domain subjects (i.e., natural science, language science, and social science), 26 topics, 127 categories, and 379 skills. Each question is annotated with explanations linked to relevant lectures from elementary and high school science curricula. We report results on the test set which includes 4,241 examples.
@@
  • POPE [37] is a benchmark for evaluating object hallucinations in MLLM generations. It includes several subsets, namely random, popular, and adversarial, which are generated using various sampling methodologies. The dataset consists of 8,910 binary classification queries, enabling thorough analysis of the object hallucination phenomena in MLLMs.
@@
  • MMBench (MMB) [43] includes approximately 3,000 multiple-choice questions, distributed across a collection of 20 distinct domains and understanding capability. Questions are designed to assess the effectiveness of MLLMs across various task paradigms. These capabilities are systematically structured into a hierarchical taxonomy, encompassing broad categories like perception and reasoning, as well as finer-grained skills such as object localization and attribute inference.
@@
  • MMMU [70] is a challenging benchmark for multimodal models, focusing on massive multi-discipline tasks demanding college-level subject knowledge. It consists of 900 validation samples drawn from university textbooks or online courses spanning six main disciplines. Questions may include multiple images interleaved with text. Evaluation includes exact and word matching for multiple-choice and open-ended questions, and models are tested in zero or few-shot settings.
@@ 4.3 Assessing the Optimal LLM Choice
Conversely, among medium-scale models, LLaVA-1.5-7B and LLaVA-1.5-LLaMA3-8B provide stronger baselines, with LLaVA-1.5-LLaMA3-8B achieving the highest score in MME-P (1544.4). However, our LLaVA-MORE models consistently surpass these baselines, demonstrating superior performance across both VQA and MLLM benchmarks. In particular, Gemma-2-9B emerges as the best-performing model, especially excelling in VQA benchmarks. It achieves the highest scores on GQA and AI2D, significantly outperforming both baselines and other models within the LLaVA-MORE family. In MLLM benchmarks, LLaVA-MORE combined with LLaMA-3.1-8B demonstrates instead strong capabilities on the MMB dataset, showing good results in multiple-choice question settings. Notably, our models exhibit less sensitivity to object hallucinations, as seen by the results achieved on the POPE benchmark, and demonstrate superior performance on the SEED dataset, further highlighting their robustness in multimodal reasoning tasks.
@@
This trend extends to MLLM benchmarks, where LLaVA-MORE models demonstrate competitive performance in MMB and SEED, further highlighting their efficiency and strong reasoning capabilities despite their smaller size. Overall, our results emphasize that scaling up is not the only path to better performance, as architectural choices and fine-tuning strategies significantly impact model effectiveness across different benchmarks.
@@ 4.4 Changing the Visual Backbone
We then investigate which visual backbone is more promising for building MLLMs. To this end, we select the best small- and medium-scale models resulting from Table 1 and study how their performance is affected when the visual encoder is varied. In detail, we opt for Phi-4-3.8B as the small-scale LLM (i.e., LLaVA-MORE-3.8B), and Gemma-2-9B as the medium-scale one (i.e., LLaVA-MORE-9B). As per the visual backbone, in Table 2, we include four pre-trained ViT-based models, in addition to the standard CLIP used by LLaVA. All models share the ViT-L/14 architecture, and yet there are striking differences in terms of training data, input image resolutions, and on the pre-training strategies. In particular, we can delineate two pre-training paradigms: DINOv2 [48], eventually enhanced with registers [19], has been pre-trained with self-supervision and knowledge distillation on 142M images, while the other visual encoders leverage the weak supervision of noisy image-text pairs during pre-training. Specifically, CLIP [51] learns robust visual models via contrastive learning, while SigLIP [71] exploits the sigmoid loss to improve cross-modal alignment. The recent SigLIP2 [66] builds upon SigLIP by incorporating additional training objectives, including image captioning, self-distillation, and image-masked prediction.
The only exception is MME-C, where DINOv2 with registers scores the highest at the 3.8B scale, and DINOv2 is the best at the 9B scale with 334.3 points. Among the models that benefited from image-text pre-training, SigLIP shows a substantial improvement over CLIP at both scales, highlighting its effectiveness despite introducing a computational overhead—specifically, it forces the LLM to process approximately 26% more visual tokens due to the use of higher-resolution inputs. Despite its more complicated training recipe, the new SigLIP2 performs on par with the original SigLIP at the 3.8B scale but records an average 0.4% gain over SigLIP at the 9B scale. For this reason, we will always include SigLIP2 as the visual backbone in all the subsequent experiments. One key finding from our evaluation is that SigLIP-based visual backbones establish a new performance frontier for MLLMs across most benchmarks. Beyond the advantage conferred by increased resolution, we attribute much of this success to the massive billion-scale image-text pre-training enabled by the sigmoid loss of SigLIP, compared to the 400M image-text pairs seen by CLIP during pre-training.
[…]
4.5 Changing the Image Resolution
The choice of a robust visual backbone and the appropriate LLM are critical factors in enhancing model performance. However, the resolution of the input image plays a similarly important role in visual understanding. Higher image resolutions provide more fine-grained visual information, which can significantly aid the MLLM in better interpreting the content and thereby improving performance.